Find the present value of an annuity due that pays $3000 at the beginning of each quarter for the next 5 years. Assume that money is worth 6.6%, compounded quarterly. (Round your answer to the nearest cent.
Answer #1

R = $3000, n = 5 * 4 = 20 quarters, i = 6.6%/4 = 1.65% (quarterly)

PV = 3000 * 16.9174 * 1.0165

PV = $51,589.7